OpenStudy (anonymous):

Solve for X 1/3 + 3/x = 1/x

OpenStudy (anonymous):

1/3+3/x=1/x 1/3=1/x-3/x 1/3=(1-3)/x 1/3=-2/x x=-6

OpenStudy (anonymous):

3/x-1/x=-1/3 2/x=-1/3 2=-1/3x x=-6

OpenStudy (anonymous):

3/x-1/x=-1/3 2/x=-1/3 2=-1/3x x=-6